OpenStudy (hunus):

Proof of n! > n^2 by induction

OpenStudy (hunus):

The problem asks me to prove that n! is greater than n^2 for n >= 4. I'm not quite sure what to do. Can anyone help push me in the right direction?

OpenStudy (anonymous):

Well, for induction, you usually end up proving the n=1 (or in this case n=4) case first. You've got that done. Then you need to identify your indictive hypothesis: e.g. n!>2n and n>4 In class the proof might look something like this: (n+1)!=n!(n+1) from the inductive hypothesis we have n!(n+1)>2n(n+1) since n>1 we have 2n(n+1)>2n(2) and 2n(2)=2n+1 Now, we can string it all togther to get the inequality: (n+1)!=n!(n+1)>2n(n+1)>2n(2)=2n+1 (n+1)!>2n+1 In general, it's worth trying to figure out wether it 'safe' to multiply n!>2n by n+1>2 while preserving the inequality. Which is really what I wanted to do in my head. As soon as you are sure it is legitemate, you're done.

OpenStudy (agreene):

\[n! = \prod_{k=1}^{n}k\] thus: \[\prod_{k=1}^{n}k > n^2\] you can use that to expand the product and just show implicitly that n!>n^2

OpenStudy (hunus):

Any idea how to do it by induction?

OpenStudy (anonymous):

Start with: 1. n≥4 2. n!>2n and try to get to (n+1)!>2n+1 Then you need to identify your indictive hypothesis: e.g. n!>2n and n>4 In class the proof might look something like this: (n+1)!=n!(n+1) from the inductive hypothesis we have n!(n+1)>2n(n+1) since n>1 we have 2n(n+1)>2n(2) and 2n(2)=2n+1 Now, we can string it all togther to get the inequality: (n+1)!=n!(n+1)>2n(n+1)>2n(2)=2n+1 (n+1)!>2n+1 In general, it's worth trying to figure out wether it 'safe' to multiply n!>2n by n+1>2 while preserving the inequality. Which is really what I wanted to do in my head. As soon as you are sure it is legitemate, you're done.

OpenStudy (hunus):

Would it be sufficient to say \[n! > n^2\] \[n!*(n+1) > n^2*(n+1)>(n+1)^2\] so \[(n+1)! > (n+1)^2\]

OpenStudy (zarkon):

I would add... \[(n+1)!\]\[=n!(n+1)>n^2(n+1)>(n^2-1)(n+1)\]\[=(n-1)(n+1)(n+1)=(n-1)(n+1)^2>(n+1)^2\]

OpenStudy (hunus):

Thanks :D